OpenStudy (anonymous):

Solve for the indicated variable. A = 1 + prt; for r

OpenStudy (anonymous):

A=1+prt We want to solve for r, so that just means we need to get r alone in the equation. A-1=1-1+prt A-1=prt (A-1)/pt=prt/pt (A-1)/pt=r r=(A-1)/pt And that's your answer! :)